The moment of inertia as well as the frequencies of two magnets are in the ratio 3:4 the ratio of their magnetic moments is ?


1
Expert's answer
2021-03-29T09:00:17-0400
"\frac{2\pi\sqrt{\frac{I_1}{M_1B}}}{2\pi\sqrt{\frac{I_2}{M_2B}}}=\frac{4}{3}\\
\frac{\sqrt{\frac{3}{M_1}}}{\sqrt{\frac{4}{M_2}}}=\frac{4}{3}\\\frac{M_1}{M_2}=\frac{27}{64}"